OpenStudy (anonymous):

2y - 9 + y - 5 can somebody explain how to solve this?

OpenStudy (anonymous):

you cannot add a number with a variable, because we don't know what y stands for. Instead, you can add and subtract the y's with each other and add and subtract the normal numbers together like so (2y+y)-(9-5) Notice how I just rearranged the problem? The problem is the same, but it's in better order. Wanna try adding the like terms now?

OpenStudy (anonymous):

2y - 4? so Y =2?

OpenStudy (anonymous):

woh no lol its multiple choice its not 2 0.o A. 9 B.11 C.13 D.27

OpenStudy (anonymous):

2y + y = 3y you forgot to add the y's it was almost my fault to put a parenthesis around the normal numbers, because 9 is negative. Since both 9 AND 5 are negative, you add them. Just adding the y's, the question is now 3y-9-5 try that now

OpenStudy (anonymous):

It was my fault* sorry, not almost. It was*

OpenStudy (anonymous):

Where did the 3 come from ? im completely lost i can answer it now i got it but for future question how do i get that equation?

OpenStudy (anonymous):

hmmmm... those answers don't seem to add up.... Is there any additional information to the problem? Is the equation equal to anything?

OpenStudy (anonymous):

the 3 came from adding 2y and y. ONE y plus TWO y's equals THREE y's

OpenStudy (anonymous):

Let W be a point between points U and V on UV. If UV = 13, UW = 2 y – 9, and WV = y – 5, solve for y

OpenStudy (anonymous):

If UV = 13, UW = 2 y - 9 , and W V = y - 5, solve for y